ID |
Date |
Author |
Type |
Category |
Subject |
3574
|
Wed Sep 15 01:58:28 2010 |
valera | Update | PSL | FSS locking |
The RefCav is locked and aligned. I changed the fast gain sign by changing the jumper setting on the TTFSS board. The RefCav visibility is 70%. The FSS loop ugf is about 80 kHz (plot attached. there is 10 dB gain in the test point path. this is why the ugf is at 10 dB when measured using in1 and in2 spigots on the front of the board.) with FSS common gain max out at 30 dB. There is about 250 mW coming out of the laser and 1 mW going to RefCav out of the back of the PMC. So the ugf can be made higher at full power. I have not made any changes to account for the PMC pole (the FSS is after the PMC now). The FSS fast gain was also maxed out at 30 dB to account for the factor of 5 smaller PZT actuation coefficient - it used to be 16 dB according to the (previous) snap shot. The RefCav TRANS PD and camera are aligned. I tuned up the phase of the error signal by putting cables in the LO and PD paths. The maximum response of the mixer output to the fast actuator sweep of the fringe was with about 2 feet of extra cable in the PD leg.
I am leaving the FSS unlocked for the night in case it will start oscillating as the phase margin is not good at this ugf. |
Attachment 1: DSC_2510.JPG
|
|
3573
|
Wed Sep 15 01:27:52 2010 |
rana, steve, valera | Update | PSL | FSS cables connected |
- connected the TTFSS cables (FSS fast goes directly to NPRO PZT for now)
- measured the reference cavity 21.5 MHz EOM drive to be 17.8 dBm
- turned on the HV for the FSS phase correcting EOM (aka PC) drive
- connected and turned on the reference cavity temperature stabilization
- connected the RefCav TRANS PD
- fine tuned the RefCav REFL PD angle |
3572
|
Tue Sep 14 18:07:41 2010 |
Alberto | Update | Electronics | Frequency Box Documentation Added to the SVN |
I completed a LIGO document describing design, construction and characterization of the RF System for the 40m upgrade.
It is available on the SVN under https://nodus.ligo.caltech.edu:30889/svn/trunk/docs/upgrade08/RFsystem/RFsystemDocument/
It can also be found on the 40m wiki (http://lhocds.ligo-wa.caltech.edu:8000/40m/Upgrade_09/RF_System#preview), and DCC under the number T1000461. |
3571
|
Tue Sep 14 00:21:51 2010 |
rana | Omnistructure | Environment | Office area temperature change |
I changed the setpoint for the HVAC control (next to Steve) from 73F to 72F. This is to handle the temperature increase in the control room with the AC unit there turned off.
We know that the control setpoint is not linear, but I hope that it settles down after several hours. Lets wait until Tuesday evening before making another change. |
3570
|
Mon Sep 13 22:51:07 2010 |
tara,valera | Configuration | PSL | beam scan for RCAV |
On Friday, Valera and I calculated the modematching for reference cavity from AOM.
We scan the beam profile where the spot should be.
The first beam waist in the AOM is 103 um, the lens (f= 183 mm, I'm not sure if I have the focal length right) is 280 mm away.
The data is attached. The first column is marking on the rail in inches,
the second column is distance from the lens, the third and fourth column are
vertical and horizontal spot radius in micron. Note that the beam is very elliptic because of the AOM. |
Attachment 1: 2010_09_10_w.mat
|
3569
|
Mon Sep 13 21:52:53 2010 |
kiwamu | Update | Green Locking | the X end laser is ON |
I turned ON the laser at the X end station, which had been OFF for several weeks because of the crane business.
Now the green beam hits the ITMX and I got a reflection back to the end table.
This green beam will be a nice reference when we install the green periscope in the chamber.
If it's necessary, feel free to correct the alignment of the green beam during my absence. |
3568
|
Mon Sep 13 19:41:38 2010 |
rana | Update | PSL | FSS AOM alignment |
The IR sensitive Olympus 570 camera gives us a really nice view of these IR beams. Its actually a lot better than what you can get with the analog IR viewers:
|
3567
|
Mon Sep 13 12:09:42 2010 |
kiwamu | Summary | General | elog restarted |
Elog didn't respond, so I restarted it with the script.
Before killing the process somehow two elog daemons were running at the same time. I killed those two manually. |
3566
|
Mon Sep 13 11:49:34 2010 |
kiwamu | Update | PSL | mode matching from PMC to IMC |
I have been working on the mode matching lenses which are sitting after the boradband EOM.
Last Friday I checked the mode profile after the first mode matching lens (f=-150mm). The measured mode was good.
According to the calculation done by ABCD software, the waist size is supposed to be 80.9 um after that lens.
The measured waists are 80.5 um for the vertical mode and 79.4 um for the horizontal mode.
The screenshot of the ABCD's result and the plot for the mode measurement are shown below.
I didn't carefully check the mode after the last convex lens (f=200mm), but it must be already good because the beam looks having a long rayleigh range.
Now the beam is reflected back from MC1 and goes to the AP table since I coarsely aligned the beam axis to the MC.

/**** fitting result ****/
w0_v = 80.4615 +/- 0.1695 [um]
w0_h = 79.4468 +/- 0.1889 [um]
z_v = -0.115234 +/- 0.0005206 [m]
z_h = -0.109722 +/- 0.0005753 [m]
|
3565
|
Mon Sep 13 11:40:50 2010 |
Alberto | Update | Electronics | Frequency Box Documentation Added to the SVN |
I uploaded all the material about the RF frequency Generation Box into the SVN under the path:
https://nodus.ligo.caltech.edu:30889/svn/trunk/docs/upgrade08/RFsystem/frequencyGenerationBox/
I structured the directory as shown in this tree:

I'm quickly describing in a section of the Rf system upgrade document with LIGO # T1000461. |
3564
|
Mon Sep 13 10:22:48 2010 |
josephb | Update | CDS | RCG bugs/feature request wiki page |
I've started a wiki page under the Upgrade 09/New CDS section regarding known bugs and pending feature requests for the Real Time Code Generator. It can be found at http://lhocds.ligo-wa.caltech.edu:8000/40m/Bugs_and_Pending_Feature_requests_for_the_RCG. If you have any ideas to improve the RCG or encounter a bug in the code generation process (say a particular part doesn't work inside subsystems for example), please note them here.
Currently there are bugs with excitation points (they don't work inside of a subsystem block) and tags (they don't respect scope and only 1 "from" tag for each "goto" tag connected to the output of a subsystem block). |
3563
|
Mon Sep 13 02:45:59 2010 |
rana | Configuration | DMF | seisBLRMS restarts |
I restarted the seisBLRMS DMF monitor by ssh'ing into mafalda and starting up a matlab session. I also have started a StripTool session on rossa by forwarding the process from op440m.
We need to get the modern EPICS installation onto these linux machines by copying what K. Thorne has done at LLO. |
3562
|
Mon Sep 13 00:19:32 2010 |
rana | Update | | VCO Driver Output power v. slider control voltage |
I measured the RF power output of the VCO Driver box as a function of slider value. I measured using the Gigatronics Handheld power meter and connected to the AOM side of the cable after the white Pasternak DC block.
* at low power levels, I believe the waveform is too crappy to get an accurate reading - that's probably why it looks non-monotonic.
* the meter has a sticker label on it saying 'max +20 dBm'. I went above +20 dBm, but I wonder if maybe the thing isn't linear up there... |
Attachment 1: vco.png
|
|
3561
|
Sun Sep 12 23:16:52 2010 |
valera | Update | | FSS mode matching |
The attached plot shows the beam scans of the beam leaking from the back mirror of the PMC to the BS cube that first turns the S-pol beam 90 deg to the AOM and then transmits the AOM double passed and polarization rotated P-pol beam to the reference cavity. The beam from the PMC is mode matched to the AOM using a single lens f=229 mm. The ABCD file is attached. The data were taken with VCO control voltage at 5 V. We then reduced the voltage to 4 V to reduce the astigmatism. Tara has the data for the beam scan in this configuration in his notebook.
The beam from AOM is mode matched to the reference cavity using a single lens f=286.5 mm. The ABCD file is attached. |
Attachment 1: fss.pdf
|
|
Attachment 2: fssaom-abcd.tiff
|
Attachment 3: fssrc-abcd.tiff
|
3560
|
Sun Sep 12 23:02:53 2010 |
valera | Update | | PMC mode matching |
Kiwamu and I found that the first lens in the PMC mode matching telescope was mislabeled. It is supposed to be PLCX-25.4-77.3-C and was labeled as such but in fact it was PLCX-25.4-103.0-C. This is why the PMC mode matching was bad. We swapped the lens for the correct one and got the PMC visibility of 82%. The attached plot shows the beam scans before and after the PMC. The data were taken with the wrong lens. The ABCD model shown in the plot uses the lens that was there at the time - PLCX-25.4-103.0-C. The model for the PMC is just the waist of 0.371 mm at the nominal location. The snap shot of the ABCD file is attached. The calculation includes the KTP for FI and LiNb for EOM with 4 cm length. The distances are as measured on the table. |
Attachment 1: pmc.pdf
|
|
Attachment 2: pmc-abcd.tiff
|
3559
|
Sun Sep 12 22:36:03 2010 |
rana | Configuration | Computers | rossa notes |
I found Sanjit's instructions for doing the Nvidia settings too complicated and so I followed these instructions from Facebook:
http://www.facebook.com/notes/centos-howtos/installing-nvidia-display-drivers-on-centos-55/399295987425
After installations, the monitors were autodetected and the Xinerama effect is working. |
3558
|
Sat Sep 11 22:42:07 2010 |
valera | Update | PSL | PSL update |
- The PMC REFL PD was moved from the temporary location to the one called for by the PSL layout (picture attached). The leakage beams were dumped.
- The FSS reference cavity was aligned using temporary periscope and scanned using NPRO temperature sweep. The amplitude of the sweep (sine wave 0.03 Hz) was set such that the PMC control voltage was going about 100 V p-p with. With rough alignment the visibility was as high as 50% - it will be better when the cavity is locked and better aligned but not better than 80% expected from the mode astigmatism that Tara and I measured on Thursday. The astigmatism appear to come from the FSS AOM as it depends on the AOM drive. We reduced the drive control voltage from 5 V to 4V beyond that the diffraction efficiency went below 50%. The FSS REFL PD was set up for this measurement as shown in the attached picture. There is also a camera in transmission not shown in the picture. |
Attachment 1: DSC_2502.JPG
|
|
Attachment 2: DSC_2505.JPG
|
|
3557
|
Sat Sep 11 03:16:51 2010 |
rana | Configuration | Computers | rossa notes |
I wiped out the old CentOS install on rossa and installed CentOS 5.5 on there. The DVDs are on a spindle in the control room; there were 2 iso's, but I only needed the first to install most things.
It still needs to get all of the usual stuff (java, flash, nvidia) installed as well as setting up the .cshrc and the NFS mount of /cvs/cds. The userID and groupID are set to 1001 as before. Whoever
sees Sanjit first should steer him towards this elog entry.
|
3556
|
Fri Sep 10 02:35:30 2010 |
rana, kiwamu, valera, tara | Configuration | PSL | update |
Over the last couple nights we got the beam into the FSS path and all the way to the IMC and out onto the AP table.
Tara and Valera have calculated a mode-matching solution for the reference cavity. It utilizes only a single lens between the AOM and the reference cavity. Valera and Steve will move the reference cavity into place in the morning.
We noticed that the layout was too tight on the end going into the MC and so we adjusted the angle of the final zig-zag. This will put the final mode-matching lens in between the final steering mirrors (which is generally undesirable) but the lens in this case is only f=400 mm. In addition, this lens may provide some more decoupling between the steering mirrors.
The whole layout has to be a little adjusted because of a calculation mistake I made in the mode-matching. I used only the nominal focal lengths from the CVI catalog and not the effective one. For the UV-grade fused silica lenses, the effective focal length is actually 20-30% longer. Today we measured that the "f=200 mm" lens we got is actually f = 238 mm. The BK7 lenses are much closer to the nominal.
We also replaced the Klinger mount ahead of the PMC with a Polanski style so that we could get the PMC REFL beam out without hitting the mount. Valera will continue to refine this section on the weekend.
Tomorrow, we will lock the MC using feedback only to the NPRO. The 0-150 V piezo driver is on the PSL table ready for action.
I also got a LCD video monitor from Frank and hooked it up on the PSL table. If we like this kind of thing, we can get many of them. They are pretty cheap. It would be handy to have 3-4 of them on the PSL and one on every of the ISC tables. They take the standard video for input and need +12V for power. Right now the one in there is looking at the PMC transmission.
The Omnigraffle layout as of tonight is attached. |
Attachment 1: 40mPSL.pdf
|
|
3555
|
Thu Sep 9 18:53:56 2010 |
Alberto | Configuration | Electronics | Busby Box, Rai's Box, SR554 in the RF cabinet |
I stored the Busby Box, the Rai's Box and the SR554 preamp in the RF cabinet down the Y arm. |
3554
|
Thu Sep 9 16:54:56 2010 |
rana | Update | SAFETY | main laser is back ON |
I turned off the laser last night to protect against the electricians. but failed to elog it. I accept the burnt toast completely. |
3553
|
Thu Sep 9 14:42:40 2010 |
josephb | Update | CDS | Updating suspensions model |
I've been working on updating the suspensions model, incorporating Koji's refinements as well as trying to simplify the model and making it less cluttered.
This had the side benefit of making an incorrect connection obvious. I had incorrectly wired the ASCPIT input to be summed into the yaw path, and the ASCYAW input into the pitch path. This has been corrected.
I've finished a single optic, and now I am in the process of propagating the changes to all the optics, as well as cleaning up the overall diagram using Rolf's new tags, which make things much less cluttered. I've attached a screenshot of the PRM optic control model, and will be updating the Matlab web export once I've updated the full model. |
Attachment 1: SingleSUS.png
|
|
3552
|
Thu Sep 9 12:02:03 2010 |
Alberto | Update | Electronics | Frequency Generation Box - Amplitude Noise Measurements |
I measured the amplitude noise of the source outputs and the EOM outputs of the Frequency Generation box.
the setup I used is shown in this diagram:

(NB It's important that the cables from the splitter to the RF and LO inputs of the mixers are the same length).
The results of the measurements are shown in the following plot:

Considerations:
1) both Crystals (29.5MHz and 11MHz) have the same noise
2) the 55MHz source's noise is bigger than the 11 MHz (~2x): the frequency multiplication and amplification that happen before it, add extra noise
3) the noise at EOM outputs is ~2x bigger than that of the relative sources
When I have the chance, I'll plot the results of my calculations of expected noise and compare them with the measurements. |
3551
|
Thu Sep 9 11:31:54 2010 |
steve | Update | PEM | accelerometer cable connection |
I turned off/on the power to the accelerometers in order to re rout their connections. I found cable connector body-nut #3 loose to Accelerometer 2X This connector should be checked for solid performance. |
3550
|
Thu Sep 9 11:17:15 2010 |
steve | Update | SAFETY | main laser is back ON |
Quote: |
Innolight 2W 1064nm main laser was turned off for more enclosure related work.
The electrician re rooting the power- conduit to the interlock under the floor and bring it through the new concrete tunnel.
Emergency laser power shut off switch was tested at entry door 104M. It worked fine.
No laser glasses required. Laser will be turned back on around 11 am today.
|
Atm1, AC power line to enclosure and interlock was extended.
Atm2, 24V line to interlock rerouted
Atm3, job is completed
SAFETY GLASSES REQUIRED ! SAFETY GLASSES REQUIRED!
The laser is turned on. Injection current set to 0.871A.
Actually some one turned the laser off last night and did not enter it into the elog ! Burned toast award is granted !
Please come forward voluntarily. |
Attachment 1: P1060813.JPG
|
|
Attachment 2: P1060814.JPG
|
|
Attachment 3: P1060816.JPG
|
|
3549
|
Thu Sep 9 08:23:45 2010 |
steve | Update | SAFETY | main laser turned off |
Innolight 2W 1064nm main laser was turned off for more enclosure related work.
The electrician re rooting the power- conduit to the interlock under the floor and bring it through the new concrete tunnel.
Emergency laser power shut off switch was tested at entry door 104M. It worked fine.
No laser glasses required. Laser will be turned back on around 11 am today. |
3548
|
Thu Sep 9 05:58:04 2010 |
kiwamu | Update | PSL | mode matching from PMC to IMC |
I started mode matching of the beam going to IMC. The work is still going on.
According to Rana's calculation (see here), I put the first lens (f=200mm) in between two steering mirrors after PMC.
The distance from PMC to the first lens was adjusted by using a metal ruler. So I believe the accuracy is something like 1mm.
I aligned the beam path going through the broadband EOM and the mode matching lenses.
I could find the optimum position for the second lens (f=-150mm) by sliding the position of the lens and measuring the mode after it.
But the optimum position looked a bit far from the EOM. It's off by about 3-4 inch from the designed position.
Somehow I feel that the beam before the second lens goes with a smaller divergence angle than that of designed.
So tomorrow I am going to restart the work from checking the mode before the second lens.
Maybe at first I should measure the mode without going through the EOM because it changes the waist position and makes the system not straightforward. |
3547
|
Wed Sep 8 16:17:58 2010 |
steve | Update | PEM | particle counter problem |
I played with the particle counter this morning. Now it's running and reading the correct numbers on the old COCHECKLIST.adl medm screen.
However, it can not be seen with dataviewer.
Koji just rebooted the frame builder and the problem was solved. |
3546
|
Wed Sep 8 14:44:37 2010 |
josephb | Update | CDS | Updated parse_mdl_to_ipc.py and feCodeGen.pl |
I updated parse_mdl_to_ipc.py to correctly work with the 3 new cdsIPCx parts, namely cdsIPCx_PCIE (for dolphin connections), cdsIPCx_RFM (for our traditional reflected memory connections), and cdsIPCx_SHMEM (local shared memory on the computer). These parts replaced cdsIPCx awhile back (see here ).
The code now correctly counts each type independantly with regards to ipcNum (I.e. you can have ipcNum = 0 for RFM and ipcNum = 0 for SHMEM for example).
I also went in and modified a few sections of the feCodeGen.pl (in /opt/rtcds/caltech/c1/core/advLigoRTS/src/epics/util/) so as to properly assign names to matrix adl files, matrix of filter bank adl files, and filter bank adl files. |
3545
|
Wed Sep 8 11:56:24 2010 |
kiwamu | Summary | CDS | September CDS test plan |
Joe and Kiwamu
We discussed about our CDS plan for this September. The summary of the plan and "to do list" are now on the wiki page;
http://lhocds.ligo-wa.caltech.edu:8000/40m/Upgrade_09/CDS/September_CDS_plan
Basically there are three major missions that we will do in this month;
1. complete damping of the vertex suspensions
2. Preparation for Green locking
3. Development of Simulated Plants
We also try to keep updating the wiki page. |
3544
|
Wed Sep 8 11:46:53 2010 |
kiwamu | Update | PSL | update of the layout |
I put some green stuff on the layout drawing.
I continue to refine the positions of these stuff.
Notes :
1. I flipped the reference cavity. So now the cavity is sitting on the left hand side of the layout.
2. I removed the ISS stuf. We should think about where ISS should be.
Quote: |
Also, Kiwamu has modified the layout drawing to add the green PLL stuff. This has collapsed the reference cavity's wave function placing it close to its original position.
WE (maybe Valera and Steve) can now put the reference cavity back on the table.
|
|
Attachment 1: 40mPSL.png
|
|
3543
|
Wed Sep 8 09:50:13 2010 |
rana | Update | PSL | PSL update |
Also, Kiwamu has modified the layout drawing to add the green PLL stuff. This has collapsed the reference cavity's wave function placing it close to its original position.
WE (maybe Valera and Steve) can now put the reference cavity back on the table. |
3542
|
Wed Sep 8 00:01:07 2010 |
rana, valera | Update | PSL | PMC update |
We ran the cables for the PMC: The RF cable for the 35.5 MHz drive was cheap and so we swapped the 29.5 MHz cable for it.
There now remain 1 RG-174 cable to drive the FSS PC (21.5 MHz) and 3 Heliax for the Kiwamu Tri-Mod EOM (11, 29.5, and 55 MHz).
We also changed the BLACK HV drive cable for the RED one (previously used for the MZ). All HV cables MUST be RED.
The BLACK cable is now used for the PMC_REFL DC.
The Heliax cables are routed onto the table - it remains a Alberto/Kiwamu job to strain relieve them and attach them to the TriMod box and EOM in the morning.
The PMC is locked and we did some partially bootless alignment and mode-matching. It locks easily on a TEM00 mode (with very poor visibility), but the
rest of the beam train can now be aligned while Valera does the PMC matching mambo. 
|
3541
|
Tue Sep 7 23:49:08 2010 |
sanjit | Configuration | Computers | aldabella network configuration |
added name server 192.169.113.20 as the first entry in /etc/resolv.conf
changed the host IPs in /etc/hosts to 192.168.xxx.yyy
made:
127.0.0.1 localhost.localdomain localhost
::1 localhost6.localdomain6 localhos6
as the first two lines of /etc/hosts
/cvs/cds mounts
on ethernet, DNS look-up works without the explicit host definitions in /etc/hosts,
but those entries are needed for wifi only connection.
|
3540
|
Tue Sep 7 23:34:15 2010 |
Kiwamu, Sanjit | Configuration | Computers | e-log |
e-log was repeatedly hanging and several attempts to start the daemon failed.
problem was solved after clearing the (firefox) browser cache, cookie, everything!!
|
3539
|
Tue Sep 7 23:17:45 2010 |
sanjit | Configuration | Computers | rossa notes |
Quote: |
* rossa needs to be able move windows between monitors: Xinerama?
|
Xinerama support has been enabled on rossa using nvidia-settings. |
3538
|
Tue Sep 7 22:21:47 2010 |
Dmass | Configuration | elog | elog restarted |
Quote: |
elog crashed on an upload. restarted and it worked fine with the same file.
|
Again. Resubmitted an old entry with just text changes. Elog hung for 5 min +. |
3537
|
Tue Sep 7 22:21:17 2010 |
Alberto | Configuration | Computers | elog restarted |
|
3536
|
Tue Sep 7 20:44:54 2010 |
Yoichi | HowTo | COMSOL Tips | COMSOL example for calculating mechanical transfer functions |
I added COMSOL example files to the 40m svn to demonstrate how to make transfer function measurements in COMSOL.
https://nodus.ligo.caltech.edu:30889/svn/trunk/comsol/MechanicalTF/
The directory also contains an (incomplete) explanation of the method in a PDF file. |
3535
|
Tue Sep 7 15:57:07 2010 |
Alberto | Configuration | Computers | Nodus connection not working. Fixed |
[Joe, Alberto]
The Nodus connection to the Martian network stopped working after someone switched cables on the Netgear router. Apparently that router doesn't like to have the 23 and 24 ports connected at the same time.
Joe fixed the connection just freeing either the 23 or the 24 port. |
3534
|
Tue Sep 7 15:31:28 2010 |
josephb, alex | Update | CDS | Binary Output working/ IO chassis fixed |
As noted previously, we were having problems with getting multiple 32 channel binary output cards working. Alex came by and we eventually tracked the problem down to an incorrect counter in the c code. This has been fixed and checked into the CDS svn repository. I tested the actual hardware and we are in fact able to turn our test LEDs on with multiple binary output boards.
Alex and I also looked at the non-functional IO chassis (the one which wouldn't sync with the 1PPS signal and wasn't turning on when the computer turned on. We discovered one corner of the trenton board wasn't screwed down and was in fact slightly warped. I screwed it down properly, straightening the board out in the process. After this, the IO chassis worked with a host interface board to the computer and started properly. We were able to see the boards attached as well with lspci. So that chassis looks to be in working condition now.
Onwards to the RFM test. |
3533
|
Tue Sep 7 14:42:02 2010 |
Dmass | Configuration | elog | elog restarted |
elog crashed on an upload. restarted and it worked fine with the same file. |
3532
|
Tue Sep 7 13:31:49 2010 |
Alberto | Update | Electronics | RF System - Frequency Distribution Box - Priority Plan |
We need a distribution unit in the LSC rack to: 1) collect the demod signals coming from the Frequency Generation Box 2) adjust the power level 3) generate 2nd harmonics (for POP) 4) distribute the demod signals to the single demodulation boards.
The base line plan is the following:

The box can be build up gradually, but the priority goes to these parts:

I need help for this work. I know exactly how to do it, I just don't have the time to do it all by myself.
Besides the Distribution Box, the demodulation part of the upgrade would still require two steps:
1) upgrade the Band Pass Filters of the demodulation boards (I have all the parts)
2) cabling from the distribution box to the demod board (one-afternoon kind of job) |
3531
|
Tue Sep 7 10:50:53 2010 |
josephb | Configuration | Computers | rossa notes |
The controls group is user id 500 by default on most new machines. Unfortunately, the user ID used across the already existing machines is 1001. One method of doing this switch is in this elog. You can also do the change of the controls ID by becoming root and using the graphical command system-config-users. This will let you change the user ID and group ID for controls to 1001. This graphical interface also lets you change the login shell.
Unfortunately, I had some minor difficulty and I ended up removing the old controls and creating a new controls account with the correct values and using tcsh. The .cshrc file has been recreated to source cshrc.40m. The controls account now has correct permissions, although some of the preferences such as background will need to be reset.
Quote: |
**** Deleted
apps/emacs
apps/linux64/firefoxold
apps/linux64/comsol (old v. 3.5)
* running up2date on rossa
* rossa needs to be able move windows between monitors: Xinerama?
* there are permissions problems: controls on rossa can't
make and delete directories made by 'controls' elsewhere.
Some sort of user# or group issue?
|
|
3530
|
Tue Sep 7 08:56:00 2010 |
Alberto | Update | Electronics | Frequency Generation Box Assembly - Phase Noise Measurements |
Here are the results of my phase noise measurements on the 7 outputs of the Frequency Generation Box. (BIN=95L applied by DTT). See attached pdf for a higher definition picture.

The plot shows that the phase noise of the 11 MHz outputs (Source, EOM modulation signal, Demodulation signal) is as low as that of the Marconi. The Marconi is limiting my measurement's resolution.
The mode cleaner signal's oscillator (29.5 MHz output, blue trace) is higher than the 11MHz above 1KHz.
The 55MHz signals have all the same phase noise (traces overlapped), and that is higher than the 11 MHz ones from about 100Hz up. i don't know what's going on.
I need to use the spare 11MHz Wenzel crsytal to have a better reference source for the measurement. |
Attachment 2: 2010-09-03_FreqBoxPhaseNoise_AllOutputsComparison_smooth.pdf
|
|
3529
|
Mon Sep 6 22:09:11 2010 |
Alberto | Update | Electronics | RF Frequency Generation Box heat sink installed and tested |
Last week I noticed that the high power amplifiers in the Frequency Generation Box became hot after 2 hours of continuous operation with the lid of the box closed. When I measured their temperature it was 57C, and it was still slowly increasing (~< 1K/hr).
According to the data sheet, their maximum recommended temperature is 65C. Above that their performances are not guaranteed anymore.
These amplifiers aren't properly dissipating the heat they produce since they sit on a plastic surface (Teflon), and also because their wing heat dissipator can't do much when the box is closed. I had to come up with some way to take out their heat.
The solution that I used for the voltage regulators (installing them on the back panel, guaranteeing thermal conduction but electrical isolation at the same time) wouldn't be applicable to the amplifiers.
I discussed the problem with Steve and Koji and we thought of building a heat sink that would put the amplifier in direct contact with the metal walls of the box.
After that, on Friday I've got Mike of the machine shop next door to make me this kind of L-shaped copper heat sink:

On Saturday, I completely removed the wing heat dissipator, and I only installed the copper heat sink on top of the amplifier. I used thermal paste at the interface.

I turned on the power, left the lid open and monitored the temperature again. After 2 hours the temperature of the amplifier had stabilized at 47C.
Today I added the wing dissipator too, and monitored again the temperature with the lid open. then, after a few hours, I closed the the box.
I tracked the temperature of the amplifier using the temperature sensors that I installed in the box and which I have attached to the heat sink.
I connected the box temperature output to C1:IOO-MC_DRUM1. With the calibration of the channel (32250 Counts/Volt), and Caryn's calibration of the temperature sensor (~110F/Volt - see LIGO DOC # T0900287-00-R), the trend that I measured was this:

Conclusion
The heat sink is avoiding the amplifier to overheat. The temperature is now compatible with that of the other component in the box (i.e., crystal oscilaltors, frequency multiplier).
Even with the lid closed the temperature is not too high.
Two things remain untested yet:
1) effect of adding a MICA interface sheet between the heat sink and the wall of the chassis. (necessary for gorund isolation)
2) effect of having all 3 amplifiers on at the same time
I am considering opening air circulation "gills" on the side and bottom of the chassis.
Also we might leave the box open and who ever wants can re- engineer the heat sink.
For posterity.
- Ideally we would like that the heat sink had the largest section area. A brick of metal on top the amplifier would be more effective. Although it would have added several pounds to the weight of the box.
- We need these amplifiers in order to have the capability to change the modulation depth up to 0.2, at least. The Mini-Circuit ZHL-2X-S are the only one available off-the-shelf, with a sufficiently low noise figure, and sufficiently high output power. |
3528
|
Mon Sep 6 21:08:44 2010 |
rana | Update | CDS | Susension model reviewed |
We must remember that we are using the Rev.B SOS Coil Drivers and not the Rev. A. 
The main change from A->B was the addition of the extra path for the bias inputs. These inputs were previously handled by the slow EPICS system and not a part of the front end. So we used to have a separate bias screen for these than the bias which is in the front end. The slow bias is what was used for the alignment to avoid overloading the range of the main coil driver path. |
3527
|
Mon Sep 6 20:38:58 2010 |
Koji | Update | CDS | Susension model reviewed |
I have reviewed the suspension model of C1SUS and refined it.
It is comaptible to the current one but has minor additions. |
Attachment 1: suspension_model.pdf
|
|
3526
|
Mon Sep 6 10:08:10 2010 |
Alberto | Configuration | Computers | Netgear Network Switch fan broken. |
The Netgear Network Switch in the top shelf of Nodus' rack has a broken fan. It is the one interfaced to the Martian network.
The fan must have broken and it is has now started to produce a loud noise. It's like a truck was parked in the room with the engine running.
Also the other network switch, just below the Netgear, has one of its two fans broken. It is the one interfaced with the General Computer Side.
I tried to knock them to make the noise stop, but nothing happened.
We should consider trying to fix them. Although that would mean disconnecting all the computers. |
3525
|
Mon Sep 6 01:46:58 2010 |
rana | Update | PSL | PSL update of the day |
To bypass the polarizer issue, I just used cubes. One I took from the FSS-Refcav path and the other from the power control part of the old MOPA, just downstream of the MOPA's periscope.
We'll swab these out with the thin-film polarizers after we get the mounts made.
With the cubes in, I also installed the Faraday + its 1/2-wave plate. The transmission looks good and we're getting into the PMC and its flashing a TEM00 mode sometimes. I set up a signal generator to drive the SLOW actuator by 1 FSR at 0.1 Hz.
I have set up a PMC transmission camera and transPD so that its easy to align. The flashing mode already allows us to align most of the rest of the table (except FSS).
Our next step should be to run the cables for locking the PMC:
- RF cables to the PMC_REFL
- Dsub for the PMC RFPD
- HV cable between the PMC servo board and the PMC PZT (why is this not RED? we have to make sure to abide by the cabling color code).
- RF cable from 35.5 MHz Frequency Reference card to the PMC EOM via the RF summing box on the table.
On Tuesday, we need to make sure that all of our mounts' drawings are in the cue for the shop. I'll put the list of mounts onto the PSL upgrade wiki page.
We also have to come up with a plan for wiring some of the 2W NPRO's channels into the cross-connect so that we can have some laser channels recorded by EPICS. |